pKa of NH4+ is 5.62*10^-10

what is the pH of solution of 0.1M NH4Cl 40m added to 0.1M HCl 20mL?
Answer #1

pKa of NH4+ is 5.62*10^-10

0.1 M HCl 20 mL = 0.020 L * 0.10 M = 0.002 mole HCl

and

0.1 M NH4Cl 40 ml = 0.040 L * 0.1 mole / L = 0.0040 mole

total volume = (20 + 40) = 60 ml = 0.060 L

thus

concentration of HCl in the mixture = 0.002 mole / 0.060 L = 0.0333 M

HCl is strong acid and hence [HCl] = [H+]

pH = - log [H+] = - log (0.0333) = 1.477

therefore,

pH of solution = 1.477